# Plug-and-Play Dramaturge: A Divide-and-Conquer Approach for Iterative Narrative Script Refinement via Collaborative LLM Agents

### v1

@@ 7. Maintain consistency with the scene’s narrative function and context
Table 13: Prompt used for the Scene-level Review Integrator.
Suggestion Router
@@
Table 14: Prompt used for the Suggestion Router.
Appendix C Appendix C: LLM prompts for Hierarchical Coordinated Revision Stage
@@ 7. Enhances emotional resonance without overwhelming changes
Table 15: Prompt used for the Storyline Editor - Brainstormer Submodule.
Decomposer Submodule
@@ 3. A beat (key moment advancing the story)
Table 16: Prompt used for the Storyline Editor - Decomposer Submodule.
Scene Editor
@@
Table 17: Prompt used for the Scene Editor.
Dialogue Editor
@@ 5. Scene Transitions:
Table 18: Prompt used for the Dialogue Editor.
Script Description Editor
@@
Title Editor Submodule
@@ 3. Reflects the themes and tone of the screenplay
Table 19: Prompt used for the Script Description Editor - Title Editor Submodule.
Character Description Editor Submodule
@@ 5. Adds psychological depth and nuance
Table 20: Prompt used for the Script Description Editor - Character Description Editor Submodule.
Script Polisher
@@
Scene Description Polisher Submodule
@@ 6. Ensure it sets up the dialogue that follows
Table 21: Prompt used for the Script Polisher - Scene Description Polisher Submodule.
@@ 5. Logical Completeness
Table 22: Prompt used for the Script Polisher - Dialogue Polisher Submodule.
Appendix D Appendix D: Evaluation Prompts
@@
Scene-Level Comparative Evaluation
@@
Component Comparative Evaluator Submodule
@@
Final Comprehensive Evaluator Submodule
@@
Script-Level Overall Evaluation
@@
Table 23: Prompt used for the Scene-Level Comparative Evaluation Assessment - Component Comparative Evaluator Submodule.
@@ 4. Scene Presentation
Table 24: Prompt used for the Scene-Level Comparative Evaluation - Final Comprehensive Evaluator Submodule.
@@ 4. Scene Presentation(25 points max):
Table 25: Prompt used for the Script-Level Overall Evaluation.
Appendix E Appendix E: Complete Examples of the Results
@@
Original Script Example
@@
Enhanced Script Example
The following is the complete enhanced version of the same script after processing through Dramaturge. This example demonstrates the comprehensive improvements in character development, narrative structure, dialogue quality, and scene presentation achieved by our divide and conquer framework.
Table 26: An example of an original script generated by Dramatron.
@@
Table 27: An example of the original script refined by our Dramaturge.
@@
Appendix F Appendix F: Statistics of the Experimental Dataset
@@
Table 28: Dataset statistics showing script length distribution across five datasets. All lengths are measured in words.

### v2

@@ 7. Maintain consistency with the scene’s narrative function and context
Priority Scoring Policy:
- Stage 1 (Global Alignment): enforce compliance with Global Review directives, discarding conflicting
local suggestions and elevating those that explicitly fulfill GR guidance.
- Stage 2 (Impact Scope): within the remaining pool, prioritize interventions whose influence propagates
to subsequent scenes or multiple dialogue exchanges over single-line or purely stylistic tweaks.
- Stage 3 (Inspector Tier): when previous stages tie, favor Plot and Character Inspector suggestions
before Dialogue and Scene Description Inspectors because they most strongly affect causality
and behavioral authenticity.
@@
B.6. Suggestion Router
@@
Appendix C LLM prompts for Hierarchical Coordinated Revision Stage
@@ C.1.1. Brainstormer Submodule
Table 16. Prompt used for the Storyline Editor - Brainstormer Submodule.
@@ 7. Enhances emotional resonance without overwhelming changes
C.1.2. Decomposer Submodule
@@
Table 17. Prompt used for the Storyline Editor - Decomposer Submodule.
@@ 3. A beat (key moment advancing the story)
C.2. Scene Editor
@@
C.3. Dialogue Editor
@@ 5. Scene Transitions:
C.4. Script Description Editor
@@
C.4.1. Title Editor Submodule
@@
Table 20. Prompt used for the Script Description Editor - Title Editor Submodule.
@@ 3. Reflects the themes and tone of the screenplay
C.4.2. Character Description Editor Submodule
@@
Table 21. Prompt used for the Script Description Editor - Character Description Editor Submodule.
@@ 5. Adds psychological depth and nuance
C.5. Script Polisher
@@
C.5.1. Scene Description Polisher Submodule
@@
Table 22. Prompt used for the Script Polisher - Scene Description Polisher Submodule.
@@ 6. Ensure it sets up the dialogue that follows
Table 23. Prompt used for the Script Polisher - Dialogue Polisher Submodule.
@@ 5. Logical Completeness
Appendix D Evaluation Prompts
@@
D.1. Scene-Level Comparative Evaluation
@@
D.1.1. Component Comparative Evaluator Submodule
@@
D.1.2. Final Comprehensive Evaluator Submodule
@@
D.2. Script-Level Overall Evaluation
@@
Table 24. Prompt used for the Scene-Level Comparative Evaluation Assessment - Component Comparative Evaluator Submodule.
@@ 4. Scene Presentation:
Table 25. Prompt used for the Scene-Level Comparative Evaluation - Final Comprehensive Evaluator Submodule.
@@ 4. Scene Presentation
Table 26. Prompt used for the Script-Level Overall Evaluation.
@@ 4. Scene Presentation(25 points max):
Appendix E Complete Examples of the Results
@@
E.1. Original Script Example
@@
E.2. Enhanced Script Example
The following is the complete enhanced version of the same script after processing through Dramaturge. This example demonstrates the comprehensive improvements in character development, narrative structure, dialogue quality, and scene presentation achieved by our divide-and-conquer framework.
Table 27. An example of an original script generated by Dramatron.
@@ E.2. Enhanced Script Example
Table 28. An example of the original script refined by our Dramaturge.
@@
Appendix F Statistics of the Experimental Dataset
@@
Table 29. Dataset statistics showing script length distribution across five datasets. All lengths are measured in words.